# Concurrent Composition for Differentially Private Continual Mechanisms

### v3

@@ 1.2. Our contributions
(B) We prove concurrent composition theorems for continual mechanisms, where an adaptive adversary may interleave queries and updates across mechanisms based on the entire history of observed outputs. As discussed for the concurrent composition of interactive mechanisms, the concurrent setting involves additional adaptivity that is not supported by composition theorems for noninteractive mechanisms. Existing results on the concurrent composition of interactive mechanisms do not apply to the concurrent composition of continual mechanisms, since in our model the adversary adaptively selects both (a) the mechanisms receiving neighboring inputs (including their privacy parameters and neighboring definition) and (b) the neighboring update sequences they receive over time, which is not the case for interactive mechanisms. We show tight privacy bounds for the following useful forms of concurrent composition of continual mechanisms:
(1) The first shows that, when the privacy-loss parameters are fixed, the standard composition theorems for noninteractive approximate differential privacy extend to the concurrent composition of continual mechanisms. (2) We extend the classical “parallel composition theorem” for noninteractive mechanisms to the concurrent parallel composition of an unbounded number of adaptively chosen interactive mechanisms with an adaptively chosen pair of neighboring datasets. (3) We construct a counterexample demonstrating that the concurrent parallel composition of an unbounded number of continual mechanisms with adaptively chosen dataset updates is not private. Then imposing an upper bound on the sum of the privacy parameter $\delta_{i}$ of the created mechanisms, we introduce a new composition theorem for concurrent parallel composition with this upper bound. (4) As a special case, our new composition theorem extends the “parallel composition theorem” for noninteractive mechanisms to the concurrent composition of an unbounded number of adaptively chosen continual mechanisms when the composed mechanisms satisfy pure differential privacy.
(C) We formalize a set of general conditions on a continual mechanism $\mathcal{M}$ that runs multiple continual sub-mechanisms such that the privacy guarantees of $\mathcal{M}$ follow directly using the above concurrent composition theorems on the sub-mechanisms, without further privacy loss. These conditions are also reformulated as a set of properties for pseudocode describing continual mechanisms.
(D) We use our concurrent composition theorems for continual mechanisms, with the aforementioned general conditions, to give a simpler and more modular privacy analysis of a recent continual histogram mechanism [13]. In the case of approximate DP, ours is the first proof that the privacy of the mechanism holds against adaptive adversaries.
(E) We how that, for every filter $\mathrm{Filt}$, the concurrent $\mathrm{Filt}$-filter composition of continual mechanisms enjoys the same privacy guarantees as the $\mathrm{Filt}$-filter composition of non-interactive mechanisms.
@@
Towards defining privacy, we consider adaptive adversaries $\mathcal{A}$ that send messages $m_{i}$ for $i=1,2,3,\dots$ that always are parsed as pairs $m_{i}=(m_{i}[0],m_{i}[1])$. Privacy is defined with respect to a verification function $f$ that takes a sequence of such pairs $(m_{1},m_{2},\ldots,m_{t})$ and outputs either $\top$ or $\bot$. When $f(m_{1},m_{2},\ldots,m_{t})=\top$, we consider the two sequences $m[0]=(m_{1}[0],m_{2}[0],\dots,m_{t}[0])$ and $m[1]=(m_{1}[1],m_{2}[1],\dots,m_{t}[1])$ to be neighboring.
@@
To go from a verification function $f$ to a definition of privacy against adaptive adversaries, we utilize the notion of an interactive post-processing mechanism (IPM) [11], which is a stateful procedure that can stand between an interactive mechanism and an adversary, or two interactive or continual mechanisms, or even two IPMs, and interact with both of them adaptively, sending and receiving messages to the mechanisms on its left and right. Specifically, we place two IPMs between the adversary $\mathcal{A}$ and the continual mechanism $\mathcal{M}$. The first is a verifier $\mathcal{V}[f]$ that, every time it receives a new message pair $m_{k}$ from $\mathcal{A}$, applies $f$ to the entire history of message pairs $(m_{1},\ldots,m_{k})$ sent by $\mathcal{A}$, and forwards $m_{k}$ onward only if $f$ returns $\top$. Otherwise $\mathcal{V}[f]$ halts the interaction. The second IPM is the identifier $\mathcal{I}$, which starts out with a secret bit $b\in\{0,1\}$, and every time it receives a message pair $(m_{i}[0],m_{i}[1])$ from $\mathcal{V}[f]$, it forwards $m_{i}[b]$ on to the mechanism $\mathcal{M}$. When $\mathcal{M}$ returns an answer $a_{i}$, $\mathcal{I}$ and $\mathcal{V}[f]$ just forward it back to $\mathcal{A}$. See Figure 2 for an illustration of an interaction between $\mathcal{A}$ and $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}[1]\!\circ^{*}\!\mathcal{M}$ consisting of three messages from $\mathcal{A}$ to $\mathcal{V}[f]$.
@@
Let $f$ be a verification function. A continual mechanism $\mathcal{M}$ is $(\epsilon,\delta)$-DP with respect to $f$ if for every adaptive adversary $\mathcal{A}$, the view of $\mathcal{A}$ when interacting with $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}[0]\!\circ^{*}\!\mathcal{M}$ is $(\epsilon,\delta)$-indistinguishable from its view when interacting with $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}[1]\!\circ^{*}\!\mathcal{M}$.
The idea of defining security against adaptive adversaries by having the adversary send pairs of messages and trying to distinguish whether the first message in each pair or the second message in each pair is being used goes back to the notion of left-or-right indistinguishability in crytography [1]. What is different here is its combination with a verification function $f$ to enforce the neighboring condition, which is necessary to capture differential privacy.
@@
With the above definitions in hand, we can formulate and prove new concurrent composition theorems for continual mechanisms. Recall that with concurrent composition of interactive mechanisms, an adaptive adversary (or honest analyst) can interleave its queries to the different interactive mechanisms being composed; for concurrent composition of continual mechanisms, we also allow the dataset updates to be interleaved with the queries and each other. (See Figure 3.)
@@
Notably, this result allows the adversary to adaptively and concurrently choose the neighboring sequences fed to each of the $\mathcal{M}_{i}$’s based on the responses received from the other mechanisms.
@@
A key step in the proof of Theorem 1.2 is to show that for any $(\epsilon_{i},\delta_{i})$-DP continual mechanism $\mathcal{M}$, the interactive mechanism $\mathcal{M}^{\prime}(\cdot)=\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(\cdot)\!\circ^{*}\!\mathcal{M}$ can be simulated by post-processing the randomized response mechanism $\mathrm{RR}_{\epsilon_{i},\delta_{i}}(\cdot)$. Lyu [19] proves such a result under the assumptions that the answer space of $\mathcal{M}$ is finite and that there exists a predetermined upper bound $T$ on the number of interactions. We remove the latter assumption and relax the former to having an infinite (but discrete) answer space.
While we focus on composing approximate DP with fixed privacy loss parameters $(\epsilon_{i},\delta_{i})$ in Theorem 1.2, we expect that there should also be analogues of it for other settings in which concurrent composition of interactive mechanisms have been studied, such as Rényi DP [19], $f$-DP [24], and adaptively chosen privacy-loss parameters [11].
@@
Next, we explore the concurrent and continual analogue of parallel composition of differential privacy. To review parallel composition, suppose that $\mathcal{M}_{1},\ldots,\mathcal{M}_{\ell}$ are each $(\epsilon_{0},\delta_{0})$-DP noninteractive mechanisms with respect to $\sim$, where each $\mathcal{M}_{i}$ expects a dataset $x_{i}$ as input, and for a parameter $k\in\{0,1,\ldots,\ell\}$, we define a neighboring relation $\sim_{k}$ on $\ell$-tuples as follows: $(x_{1}[0],\ldots,x_{\ell}[0])\sim_{k}(x_{1}[1],\ldots,x_{\ell}[1])$ if:
  1. (1)
    For every $i\in[\ell]$, $x_{i}[0]\sim x_{i}[1]$ or $x_{i}[0]=x_{i}[1]$.
  2. (2)
    For at most $k$ values of $i\in[\ell]$, $x_{i}[0]\neq x_{i}[1]$.
We refer to the mechanism $\mathcal{M}^{\prime}(\cdot)=(\mathcal{M}_{1}(\cdot),\ldots,\mathcal{M}_{\ell}(\cdot))$ with neighbor relation $\sim_{k}$ as the $k$-sparse parallel composition of $\mathcal{M}_{1},\ldots,\mathcal{M}_{\ell}$. If $\mathcal{M}_{1},\ldots,\mathcal{M}_{\ell}$ are noninteractive mechanism, then it is known that $\mathcal{M}^{\prime}$ satisfies the same level of $(\epsilon,\delta)$-DP as the (standard) composition of $k$ (rather than $\ell$) $(\epsilon_{0},\delta_{0})$-DP mechanisms.
We study the extent to which parallel composition extends to the concurrent composition of continual mechanisms, meaning the neighboring relation is as above, but the adversary can interact concurrently with the mechanisms, and in particular decide adaptively on which $k$ of them to have different update sequences. Moreover, we also consider the case where $\ell$ is unbounded, meaning the adversary (or honest analyst) can adaptively decide how many mechanisms to interact with. We refer to this as concurrent $k$-sparse parallel composition. A more restricted version of this composition, which is limited to noninteractive mechanisms, was previously studied in [20].
First, we extend the parallel composition of noninteractive mechanisms to concurrent composition of interactive mechanisms (where there are adaptive queries but datasets $x_{i}$ are static, rather than dynamically updated).
@@
As discussed later in the paper, we also prove a generalization of Theorem 1.3 for the case of composing interactive mechanisms $\mathcal{M}_{i}$ with differing $(\epsilon_{i},\delta_{i})$ parameters.
Then we turn to the case of continual mechanisms, where there can also be adaptive dataset updates. We prove a negative result for approximate DP ($\delta_{0}>0$):
@@
This theorem says that, at least in terms of the $\delta$ parameter, concurrent parallel composition of continual mechanisms does not provide any benefit over non-parallel composition, since the bound of $1-(1-\delta_{0})^{\ell}$, is the bound that we would get by just applying composition over all $\ell$ mechanisms. Inspired by this counterexample, we prove a weaker version of Theorem 1.3 for the concurrent parallel composition of continual mechanisms.
@@
For every $\epsilon_{0}>0$, $0\leq\delta^{\prime}\leq 1$, and $k\in\mathbb{N}$, the concurrent $k$-sparse parallel composition of $(\epsilon_{0},\delta_{i})$-DP continual mechanisms, where the parameters $\delta_{i}$ are chosen adaptively such that $\sum_{i}\delta_{i}\leq\delta^{\prime}$, satisfies $(\epsilon,\delta+\delta^{\prime})$-DP if the composition of $k$ noninteractive $(\epsilon_{0},0)$-DP mechanisms is $(\epsilon,\delta)$-DP.
We note that the formal version of this result applies to $(\epsilon_{i},\delta_{i})$-DP continual mechanisms with possibly different $(\epsilon_{i},\delta_{i})$ parameters. The proof of Theorem 1.5 relies on a key property of our new post-processing construction (recall that $\mathrm{RR}_{\epsilon,\delta}(b)$ denotes the randomized response mechanism with parameters $\epsilon$ and $\delta$ and input $b\in\{0,1\}$):
Lemma 1.6 (post-processing of randomized response mechanisms, informal statement).
@@
Setting $\delta^{\prime}=0$, Theorem 1.5 implies:
@@
For every $\epsilon_{0}>0$ and integer $k\in\mathbb{N}$, the concurrent $k$-sparse parallel composition of $(\epsilon_{0},0)$-DP mechanisms satisfies the same $(\epsilon,\delta)$-DP guarantees as the composition of $k$ noninteractive $(\epsilon_{0},0)$-DP mechanisms.
Like Theorem 1.5, we also prove a generalization of Corollary 1.7 for the case of composing mechanisms $\mathcal{M}_{i}$ with differing $\epsilon_{i}$ parameters.
@@
Let $\mathrm{Filt}:\left([0,\infty)\times[0,1]\right)^{*}\to\{\top,\bot\}$ be a filter. The concurrent $\mathrm{Filt}$-filter composition of continual mechanisms satisfies the same $(\epsilon,\delta)$-DP guarantee as the $\mathrm{Filt}$-filter composition of non-interactive mechanisms.
@@
Composition theorems for differential privacy have two main uses. One is to bound the accumulated privacy loss when DP systems are used repeatedly over time. Another is to facilitate the design of complex DP algorithms from simpler ones. The concurrent composition theorems for interactive mechanisms have already found an application in the privacy analysis of matching algorithms in local edge differential privacy model [4]. (This work also has a result in the continual observation model, but the privacy is only proven for oblivious adversaries, which allows it to be proven without using a composition theorem for continual mechanisms.) Our concurrent composition theorem allows for a simpler and more modular privacy analysis of continual mechanisms built out of other continual mechanisms.
There are various applications, where a continual mechanism executes multiple differentially private mechanisms, such as the Sparse Vector Technique (SVT) and continual counters, as sub-mechanisms and concurrently interacts with them. Let $\mathcal{M}$ denote such a composite mechanism. Upon receiving an input, $\mathcal{M}$ may create new sub-mechanisms, interact multiple times with the existing ones, and then return an output. $\mathcal{M}$ generates inputs for its sub-mechanisms based on its own inputs and the previous outputs of all sub-mechanisms.(See Figure 4 for an illustration.) Analyzing the privacy of such mechanism designs is challenging because the creation and inputs of sub-mechanisms are interdependent and adaptive. Even if the inputs of $\mathcal{M}$ are non-adaptive, the inputs to its sub-mechanisms may become adaptive: Let $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ denote the sub-mechanisms of $\mathcal{M}$. If the inputs of $\mathcal{M}_{1}$ depend on the outputs of $\mathcal{M}_{2}$ and the inputs of $\mathcal{M}_{2}$ depend on the outputs of $\mathcal{M}_{1}$, then $\mathcal{M}_{1}$ receives inputs correlated with its own past outputs. Later, we see an example of this situation, where a histogram mechanism creates $d$ private continual counters at the initialization and an adaptive number of SVTs and Laplace mechanisms over time. In that design, the inputs of the counters depend on the output of the (active instance of the) SVT and vice versa.
Continual mechanisms are typically presented using pseudocode (rather than as formal state machines exchanging messages, which are used by our conditions on $\mathcal{M}$), where the pseudocode receives an input, updates its variables, and returns an output. We therefore reformulate our conditions in this setting. We call a variable a decision variable if it is either the output of a private sub-mechanism or computed solely based on other decision variables. Our conditions are now as follows:
  • Destination condition: Decisions about whether to produce an output, instantiate a new sub-mechanism (and of which type), or interact with an existing one (and which one) must be based only on decision variables.
  • Response condition: The value of $\mathcal{M}$’s output must be determined solely by decision variables.
  • Mapping condition: Fixing the values of decision variables, neighboring inputs to $\mathcal{M}$ must be mapped to input sequences for the sub-mechanisms with a structure for which concurrent composition privacy is private (This condition is referred to as $f\to f^{\prime}$ property in Section 6).

### v4

@@ 1. Introduction
(C) We formalize a set of general conditions on a continual mechanism $\mathcal{M}$ that runs multiple continual sub-mechanisms such that the privacy guarantees of $\mathcal{M}$ follow directly by using the above concurrent composition theorems on the sub-mechanisms, without further privacy loss (Theorem 6.1). These conditions are also reformulated as a set of properties for pseudocode describing continual mechanisms (Corollary 6.6).
(D) We use our concurrent composition theorems for continual mechanisms, with the aforementioned general conditions, to give a simpler and more modular privacy analysis of a recent continual histogram mechanism (henzingerSS2023, ). In the case of approximate DP, ours is the first proof that the privacy of the mechanism holds against adaptive adversaries.
(E) We model local protocols in the local differential privacy model as continual mechanisms with a modular structure, and demonstrate how our concurrent composition theorems for continual mechanisms—particularly the concurrent parallel composition theorem for interactive mechanisms—can simplify their privacy analysis (Theorem 8.7 and Corollary 8.8). As an example, we give a simple privacy analysis for the core decomposition local protocol proposed by Dhulipala, Henzinger, Li, Liu, Sricharan, and Zhu (dhulipala2025near, ).
@@
Towards defining privacy, we consider adaptive adversaries $\mathcal{A}$ that send messages $m_{i}$ for $i=1,2,3,\dots$ that are always parsed as pairs $m_{i}=(m_{i}[0],m_{i}[1])$. Privacy is defined with respect to a verification function $f$ that takes a sequence of such pairs $(m_{1},m_{2},\ldots,m_{t})$ and outputs either $\top$ or $\bot$. When $f(m_{1},m_{2},\ldots,m_{t})=\top$, we consider the two sequences $m[0]=(m_{1}[0],m_{2}[0],\dots,m_{t}[0])$ and $m[1]=(m_{1}[1],m_{2}[1],\dots,m_{t}[1])$ to be neighboring.
@@
To go from a verification function $f$ to a definition of privacy against adaptive adversaries, we utilize the notion of an interactive post-processing mechanism (IPM) (haney2023concurrent, ), which is a stateful procedure that can stand between an interactive mechanism and an adversary, or two interactive or continual mechanisms, or even two IPMs, and interact with both of them adaptively, sending and receiving messages to the mechanisms on its left and right. Specifically, as illustrated in Figure 1, we place two IPMs between the adversary $\mathcal{A}$ and the continual mechanism $\mathcal{M}$. The first is a verifier $\mathcal{V}[f]$ that, every time it receives a new message pair $m_{k}$ from $\mathcal{A}$, applies $f$ to the entire history of message pairs $(m_{1},\ldots,m_{k})$ sent by $\mathcal{A}$, and forwards $m_{k}$ onward only if $f$ returns $\top$. Otherwise $\mathcal{V}[f]$ halts the interaction. The second IPM is the identifier $\mathcal{I}$, which starts out with a secret bit $b\in\{0,1\}$, and every time it receives a message pair $(m_{i}[0],m_{i}[1])$ from $\mathcal{V}[f]$, it forwards $m_{i}[b]$ on to the mechanism $\mathcal{M}$. The identifier $\mathcal{I}$ with bit $b$ is shown by $\mathcal{I}(b)$. When $\mathcal{M}$ returns an answer $a_{i}$, $\mathcal{I}$ and $\mathcal{V}[f]$ just forward it back to $\mathcal{A}$. We denote the interactive mechanism representing the combination of $\mathcal{M}$, $I$, and $\mathcal{V}[f]$ by $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}\!\circ^{*}\!\mathcal{M}$.
@@
Illustration of the interactions between $\mathcal${A}\$, $\mathcal${V}[f]\$, $\mathcal{I}$(b)\$, and $\mathcal{M}$$
@@
Let $f$ be a verification function. A continual mechanism $\mathcal{M}$ is $(\epsilon,\delta)$-DP with respect to $f$ if for every adaptive adversary $\mathcal{A}$, the view of $\mathcal{A}$ when interacting with $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!\mathcal{M}$ is $(\epsilon,\delta)$-indistinguishable from its view when interacting with $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(1)\!\circ^{*}\!\mathcal{M}$.
The idea of defining security against adaptive adversaries by having the adversary send pairs of messages and trying to distinguish whether the first message in each pair or the second message in each pair is being used goes back to the notion of left-or-right indistinguishability in cryptography (BellareDJR97, ). What is different here is its combination with a verification function $f$ to enforce the neighboring condition, which is necessary to capture differential privacy.
@@
With the above definitions in hand, we can formulate and prove new concurrent composition theorems for continual mechanisms. Recall that with concurrent composition of interactive mechanisms, an adaptive adversary (or honest analyst) can interleave its queries to the different interactive mechanisms being composed; for concurrent composition of continual mechanisms, we also allow the dataset updates to be interleaved with the queries and each other. See Figure 2 for illustration.
[…]
Illustration of the interactions between an honest adversary $\mathcal${A}\$ and the concurrent composition of \$k\$ continual mechanisms
@@
Notably, this result allows the adversary to adaptively and concurrently choose the neighboring sequences fed to each of the $\mathcal{M}_{i}$’s based on the responses received from the other mechanisms. Analogous results are shown for Rényi DP in Corollary 13.4 and, under the finite-communication assumption, for $f$-DP in Corollary 14.7.
@@
A key step in the proof of Theorem 1.2 is to show that for any $(\epsilon_{i},\delta_{i})$-DP continual mechanism $\mathcal{M}$, the interactive mechanism $\mathcal{M}^{\prime}(\cdot)=\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(\cdot)\!\circ^{*}\!\mathcal{M}$ can be simulated by post-processing the randomized response mechanism $\mathrm{RR}_{\epsilon_{i},\delta_{i}}(\cdot)$. Lyu (lyu2022composition, ) proves such a result under the assumptions that the answer space of $\mathcal{M}$ is finite and that there exists a predetermined upper bound $T$ on the number of interactions. Vadhan and Zhang (vadhan2022concurrent, ) additionally assume that the query set is finite. We remove the latter assumption and relax the former to having an infinite (but discrete) answer space.
@@
Next, we explore the concurrent and continual analogue of parallel composition of differential privacy. To review parallel composition, suppose that $\mathcal{M}_{1},\ldots,\mathcal{M}_{\ell}$ are each $(\epsilon_{0},\delta_{0})$-DP noninteractive mechanisms with respect to a neighbor relation $\sim^{*}$, where each $\mathcal{M}_{i}$ expects a dataset $x_{i}$ as input, and for a parameter $k\in\{1,\ldots,\ell\}$, we define a neighboring relation $\sim^{(k)}$ on $\ell$-tuples as follows: $(x_{1},\ldots,x_{\ell})\sim^{(k)}(x_{1}^{\prime},\ldots,x_{\ell}^{\prime})$ if and only if $x_{i}=x_{i}^{\prime}$ for all but at most $k$ indices $i\in[\ell]$, for which $x_{i}\sim^{*}x_{i}^{\prime}$. We refer to the mechanism $\mathcal{M}^{\prime}(\cdot)=(\mathcal{M}_{1}(\cdot),\ldots,\mathcal{M}_{\ell}(\cdot))$ with neighbor relation $\sim^{(k)}$ as the $k$-sparse parallel composition of $\mathcal{M}_{1},\ldots,\mathcal{M}_{\ell}$. If $\mathcal{M}_{1},\ldots,\mathcal{M}_{\ell}$ are noninteractive mechanisms, then it is known that $\mathcal{M}^{\prime}$ satisfies the same level of $(\epsilon,\delta)$-DP as the (standard) composition of $k$ (rather than $\ell$) $(\epsilon_{0},\delta_{0})$-DP mechanisms.
We extend this notion to the concurrent $k$-sparse parallel composition of continual mechanisms, where an adversary (or an honest analyst) may adaptively create continual mechanisms (with no bound $\ell$ on the number of mechanism), where each mechanism satisfies $(\epsilon_{0},\delta_{0})$-DP with respect to a verification function $f^{*}$, and interact with them concurrently. Formally, instead of fixing one dataset pair $(x_{i},x_{i}^{\prime})$ for each mechanism in advance, the adversary issues pairs of messages $m=(m^{0},m^{1})$ to mechanisms over time. All but at most $k$ mechanisms receive identical message pairs throughout the interaction, while for each of the $k$ exceptional mechanisms, the sequence of message pairs $\sigma$ satisfies $f^{*}(\sigma)=\top$. If messages are data records to be inserted in the dataset, then this concurrent composition captures adaptive partitioning of data among mechanisms over time. Note that an adversary can decide to designate a mechanism $\mathcal{M}_{i}$ as one of the $k$ “exceptional” mechanisms, i.e., issue at least one pair of non-identical messages to $\mathcal{M}_{i}$, after potentially interacting with other mechanisms and sending multiple identical message pairs to $\mathcal{M}_{i}$ itself.
First, we extend the parallel composition of noninteractive mechanisms to concurrent composition of interactive mechanisms (where there are adaptive queries but datasets $x_{i}$ are static, rather than dynamically updated). [^8] As an intermediate step, Qiu et al.(qiu2022differential, ) analyzes the composition of infinitely many SVT instances (sharing a secret dataset) that are executed sequentially, where each instance begins only after the previous one halts, and where a differing element between two neighboring datasets can affect the output of at most one instance. Their result is weaker than ours as it does not allow for concurrent interaction of the adversary with the interactive mechanisms. We recall that an interactive mechanism is modeled as a continual mechanism receiving a dataset as its first message and queries as subsequent messages. (The verification function $f^{*}$ for this mechanism ensures that the first message is a pair of neighbor datasets, while the subsequent pairs are identical queries.) A formal statement of the following theorem is given in Corollary 5.15, which more generally composes interactive mechanisms $\mathcal{M}_{i}$ with differing $(\epsilon_{i},\delta_{i})$ parameters.
@@
We prove analogous results for Rényi DP in Corollary 13.7 and, under the finite-communication assumption, for $f$-DP in Corollary 14.9. Then we turn to the case of continual mechanisms, where there can also be adaptive dataset updates. We prove a negative result for approximate DP ($\delta_{0}>0$):
@@
This theorem says that, at least in terms of the $\delta$ parameter, concurrent parallel composition of continual mechanisms does not provide any benefit over non-parallel composition, since the bound of $1-(1-\delta_{0})^{\ell}$ is the bound that we would get by just applying composition over all $\ell$ mechanisms. The high-level idea of this counterexample is to define a continual mechanism $\mathcal{M}_{0}$ that, upon receiving its first input message (regardless of the value), outputs $\bot$ with probability $\delta$ and $\top$ otherwise. If the first output is $\bot$, then on its second input the mechanism simply returns that input, revealing it entirely. Consider an adversary $\mathcal{A}$ that creates $\ell$ instances of $\mathcal{M}_{0}$, sends the pair $(0,0)$ to each instance, and then sends the pair $(0,1)$ to the first mechanism whose response was $\bot$. With probability $1-(1-\delta)^{\ell}$, such a mechanism exists and its second response reveals its input bit, which equals the secret bit used by the identifier to filter the adversary’s message pairs. An analogous counterexample for Rényi DP is provided in Theorem 13.5.
Inspired by this counterexample, we prove a weaker version of Theorem 1.3 for the concurrent parallel composition of continual mechanisms, stated below. The formal version of this result is given in Theorem 5.4 and applies to $(\epsilon_{i},\delta_{i})$-DP continual mechanisms with possibly different $(\epsilon_{i},\delta_{i})$ parameters.
@@
For every $\epsilon_{0}>0$, $0\leq\delta^{\prime}\leq 1$, and $k\in\mathbb{N}$, the concurrent $k$-sparse parallel composition of $(\epsilon_{0},\delta_{i})$-DP continual mechanisms, where the parameters $\delta_{i}$ are chosen adaptively such that $\prod_{i}(1-\delta_{i})\geq 1-\delta^{\prime}$, satisfies $(\epsilon,1-(1-\delta)(1-\delta^{\prime}))$-DP if the composition of $k$ noninteractive $(\epsilon_{0},0)$-DP mechanisms is $(\epsilon,\delta)$-DP.
The proof of Theorem 1.5 relies on a key property of our new post-processing construction (recall that $\mathrm{RR}_{\epsilon,\delta}(b)$ denotes the randomized response mechanism with parameters $\epsilon$ and $\delta$ and input $b\in\{0,1\}$):
Proposition 1.6 (post-processing of randomized response mechanisms, informal statement).
@@
The first sentence of this proposition was proven by Lyu (lyu2022composition, ) and Vadhan and Zhang (vadhan2022concurrent, ) with finiteness condition on the communication. The novelty in our result is that $\mathcal{P}$ does not access the outcome of $\mathrm{RR}_{\epsilon,0}(b)$ until $\mathcal{M}(x_{0})$ and $\mathcal{M}(x_{1})$ differ. Our construction builds on Lyu’s proof, where an IPM is initially given the outcome of $\mathrm{RR}_{\epsilon,\delta}(b)$. Let $\mathcal{M}^{\prime}$ be a continual mechanism that is $(\epsilon,\delta)$-DP with respect to a verification function $f$. Consider the interaction between $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}\!\circ^{*}\!\mathcal{M}^{\prime}$ and an adversary. As long as the adversary sends pairs of identical messages, the answer distributions of $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!\mathcal{M}^{\prime}$ and $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(1)\!\circ^{*}\!\mathcal{M}^{\prime}$ will be identical. Thus, by replacing $\mathcal{M}(x_{b})$ in Proposition 1.6 with $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!\mathcal{M}^{\prime}$, we conclude that there exists an IPM $\mathcal{P}$ that simulates $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!\mathcal{M}^{\prime}$ using only the output of $\mathrm{RR}_{0,\delta}(b)$ until the adversary issues a pair of non-identical messages. At that point, $\mathcal{P}$ requires the output of $\mathrm{RR}_{\epsilon,0}(b)$ to continue its simulation. This result is the core idea to show that in the concurrent $k$-sparse parallel composition of continual mechanisms, the $\epsilon$ parameters of the (potentially unboundedly many) composed mechanisms that are always assigned pairs of identical messages do not contribute to the overall privacy loss.
In the concurrent parallel composition of interactive mechanisms, the adversary sends to each mechanism a pair of identical or neighboring datasets as first message and identical query pairs as subsequent messages. The key idea in the proof of Theorem 1.3 is that we can determine whether a mechanism is one of the $k$ exceptional ones solely from its first pair, before the mechanism produces any output. Recall that all messages of the adversary are filtered by a secret bit $b\in\{0,1\}$. If a mechanism is non-exceptional, whether the bit $b$ is zero or one makes no difference, and the first element of every pair can always be forwarded. In this proof, we simulate a mechanism by post-processing of a randomized-response mechanism with input $b$ only if it is exceptional, and show that we incur no privacy cost—neither in $\epsilon$ nor in $\delta$—from the non-exceptional mechanisms.
Setting $\delta^{\prime}=0$ in Theorem 1.5 yields the following corollary, stated formally (and for differing $\epsilon_{i}$) in Corollary 1.7.
@@
For every $\epsilon_{0}>0$ and integer $k\in\mathbb{N}$, the concurrent $k$-sparse parallel composition of $(\epsilon_{0},0)$-DP continual mechanisms satisfies the same $(\epsilon,\delta)$-DP guarantees as the composition of $k$ noninteractive $(\epsilon_{0},0)$-DP mechanisms. [^9] In some composition theorems (e.g., advanced composition), composing pure DP mechanisms yields $(\epsilon,\delta)$-DP with $\delta>0$.
@@
Let $\mathrm{Filt}:\left([0,\infty)\times[0,1]\right)^{*}\to\{\top,\bot\}$ be a filter. The concurrent $\mathrm{Filt}$-filter composition of continual mechanisms satisfies the same $(\epsilon,\delta)$-DP guarantees as the $\mathrm{Filt}$-filter composition of noninteractive mechanisms.
This result is also shown for Rényi DP in Corollary 13.8 and for $f$-DP, under the finite-communication assumption, in Corollary 14.10.
@@
Composition theorems for differential privacy have two main uses. One is to bound the accumulated privacy loss when DP systems are used repeatedly over time. Another is to facilitate the design of complex DP algorithms from simpler ones, for example in modular software designs. The concurrent composition theorems for interactive mechanisms have already found an application in the privacy analysis of matching algorithms in local edge differential privacy model (DinitzLiLiZh25, ). (That work also has a result in the continual observation model, but the privacy is only proven for oblivious adversaries, which allows it to be proven without using a composition theorem for continual mechanisms.) Our concurrent composition theorem allows for a simpler and more modular privacy analysis of continual mechanisms built out of other continual mechanisms.
There are various applications where a continual mechanism executes multiple differentially private mechanisms, such as the Sparse Vector Technique (SVT) and continual counters, as sub-mechanisms and concurrently interacts with them. Let $\mathcal{M}$ denote such a composite mechanism. Upon receiving an input, $\mathcal{M}$ may create new sub-mechanisms, interact multiple times with the existing ones, and then return an output. $\mathcal{M}$ generates inputs for its sub-mechanisms based on its own inputs and the previous outputs of all sub-mechanisms. (See Figure 3 for an illustration.) Analyzing the privacy of such mechanism designs is challenging because the creation and inputs of sub-mechanisms are interdependent and adaptive. Even if the inputs of $\mathcal{M}$ are non-adaptive, the inputs to its sub-mechanisms may become adaptive: Let $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ denote the sub-mechanisms of $\mathcal{M}$. If the inputs of $\mathcal{M}_{1}$ depend on the outputs of $\mathcal{M}_{2}$ and the inputs of $\mathcal{M}_{2}$ depend on the outputs of $\mathcal{M}_{1}$, then $\mathcal{M}_{1}$ receives inputs correlated with its own past outputs. Later, we see an example of this situation, where a histogram mechanism creates $d$ private continual counters at the initialization and an adaptive number of SVTs and Laplace mechanisms over time. In that design, the inputs of the counters depend on the output of the (active instance of the) SVT and vice versa.
@@
Illustration of the interactions between an honest adversary $\mathcal${A}\$ and a mechanism with \$k\$ sub-mechanisms
Continual mechanisms are typically presented using pseudocode (rather than as formal state machines exchanging messages, which are used by our conditions on $\mathcal{M}$), where the pseudocode receives an input, updates its variables, and returns an output. We therefore reformulate our conditions in this setting. We call a variable an untainted variable if it is either the output of a private sub-mechanism or computed solely based on other untainted variables. Our conditions are now as follows:
  • Destination condition: Decisions about whether to produce an output, instantiate a new sub-mechanism (and of which type), or interact with an existing one (and which one) must be based only on untainted variables.
  • Response condition: The values of $\mathcal{M}$’s outputs must be determined solely by untainted variables.
  • Mapping condition: Fixing the values of untainted variables, neighboring inputs to $\mathcal{M}$ must be mapped to input sequences for the sub-mechanisms with a structure for which concurrent composition is private (This condition is referred to as the $f\to f^{\prime}$ property in Section 6.)